# Robustness of topological phases on aperiodic lattices

### v1

@@ 3.2.2. Position spectral triple generates K-homology
We claim that the following diagram commutes:
@@
Here the horizontal arrows are induced by inclusion maps
$\mathrm{C}_{0}(\mathbb{R}^{0,d})\hookrightarrow\mathrm{C}(\mathbb{T}^{d}),\qquad\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}\hookrightarrow\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})_{\mathbb{R}}.$
The first map comes from an inclusion of “real” manifolds $\mathbb{R}^{0,d}\hookrightarrow\mathbb{T}^{d}$. The map $\alpha_{\mathbb{R}^{0,d}}$ is the generator of the Bott periodicity of the “real” manifold $\mathbb{R}^{0,d}$, mapping a generator of $\operatorname{KKR}(\mathbb{C},\mathrm{C}_{0}(\mathbb{R}^{0,d})\otimes\mathbb{C}\ell_{0,d})$ to a generator of $\operatorname{KO}_{0}(\mathbb{R})\simeq\mathbb{Z}$, which is unique up to a sign. The vertical isomorphism $\operatorname{KKR}(\mathbb{C},\mathrm{C}(\mathbb{T}^{d})\otimes\mathbb{C}\ell_{0,d})$ is the composition
$\operatorname{KKR}(\mathbb{C},\mathrm{C}(\mathbb{T}^{d})\otimes\mathbb{C}\ell_{0,d})\xrightarrow{\operatorname{Ad}_{U^{*}}}\operatorname{KKR}(\mathbb{C},\mathrm{C}^{*}(\mathbb{Z}^{d})\otimes\mathbb{C}\ell_{0,d})\\ \xrightarrow{\sim}\operatorname{KKO}(\mathbb{R},\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\mathrm{C}\ell_{0,d})\xrightarrow{\sim}\operatorname{KKO}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}).$
The map $\alpha_{\mathbb{T}^{d}}$ is given by Kasparov’s Dirac element of the “real” manifold $(\mathbb{T}^{d},\mathfrak{r}\colon z\mapsto\overline{z})$, represented by the the Hodge–de Rham operator $\mathrm{d}+\mathrm{d}^{*}$. It follows from Lemma 2.8 that, via the Fourier transform isomorphism (2.7), the Hodge–de Rham operator $\mathrm{d}+\mathrm{d}^{*}$ is mapped to the operator ${\scriptstyle\mathsf{X}}=\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}$. We may further replace $\operatorname{KKR}(\mathrm{C}^{*}(\mathbb{Z}^{d})\otimes\mathbb{C}\ell_{d,0},\mathbb{C})$ by $\operatorname{KKO}(\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\mathrm{C}\ell_{d,0},\mathbb{R})$. This maps $\alpha_{\mathbb{T}^{d}}$ to the class represented by the real spectral triple
(3.29) $\left(\mathcal{A}\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right),$
where $\mathcal{A}\subseteq\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}$ is a dense, *-subalgebra. It is the pullback of the class of $\xi^{\mathrm{Roe}}_{\mathbb{Z}^{d}}$ in $\operatorname{KK}(\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)\otimes\mathrm{C}\ell_{0,d},\mathbb{R})$ of the spectral triple (3.27). This shows that the right part of the diagram commutes.
The left triangle of the diagram commutes, as was proven in [Ewert-Meyer:Coarse_geometry]*Proposition 11, because the inclusion $\mathbb{R}^{0,d}\hookrightarrow\mathbb{T}^{d}$ preserves the involution $\tau\colon z\mapsto-z$ of $\mathbb{T}^{d}$, thus gives an ideal inclusion $\mathrm{C}_{0}(\mathbb{R}^{0,d},\mathbb{C}\ell(\mathbb{R}^{0,d}))\hookrightarrow\mathrm{C}(\mathbb{T}^{d},\mathbb{C}\ell(\mathbb{T}^{d}))$. This ideal inclusion maps the Dirac element of $\mathbb{R}^{0,d}$ to the Dirac element of $\mathbb{T}^{d}$ via the Kasparov product.
Finally, both $\operatorname{KKR}(\mathbb{C},\mathrm{C}_{0}(\mathbb{R}^{0,d})\otimes\mathbb{C}\ell_{0,d})$ and $\operatorname{KKO}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})_{\mathbb{R}})$ are isomorphic to the abelian group $\mathbb{Z}$. Thus the commutativity of the diagram claims that the image of $\xi_{\mathbb{Z}^{d}}^{\mathrm{Roe}}$ must be the image of $\alpha_{\mathbb{T}^{d}}$. This gives a surjective group homomorphism $\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})_{\mathbb{R}})\to\mathbb{Z}$. Therefore, it follows from the K-theory of $\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})_{\mathbb{R}}$ in (3.21) that this map must be an isomorphism. ∎
Lemma 3.30.
Let $L_{1},L_{2}\subseteq\mathbb{R}^{d}$ be Delone sets and set $L_{0}\mathrel{\vcentcolon=}L_{1}\cup L_{2}$.
  • For $i\in\{0,1,2\}$, let $\mathcal{H}_{i}\mathrel{\vcentcolon=}\ell^{2}(L_{i},\mathcal{K})$ be the standard ample $L_{i}$-modules of the discrete space $L_{i}$; and $\xi_{i}^{\mathrm{Roe}}\mathrel{\vcentcolon=}\xi^{\mathrm{Roe}}_{L_{i}}$ for the position spectral triples over $\mathrm{C}^{*}_{\mathrm{Roe}}(L_{i})$ as in (3.25).
  • For $i\in\{1,2\}$, let $\iota_{i}\colon L_{i}\hookrightarrow L_{0}$ be the isometric embeddings, and $V_{i}\colon\mathcal{H}_{i}\to\mathcal{H}_{0}$ be the isometries between Hilbert spaces induced by $\iota_{i}$, that is, $V_{i}\lvert x\rangle\mathrel{\vcentcolon=}\lvert\iota_{i}(x)\rangle$ for all $x\in L_{i}$.
Then the followings hold:
  1. (1)
    $V_{i}^{\prime}s$ are covering isometries for $\iota_{i}$. Hence $\operatorname{Ad}_{V_{i}}\colon T\mapsto V_{i}TV_{i}^{*}$ maps $\mathrm{C}^{*}_{\mathrm{Roe}}(L_{i})$ into $\mathrm{C}^{*}_{\mathrm{Roe}}(L)$, and induce isomorphisms
    $\iota_{i*}\colon\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}_{\mathrm{Roe}}(L_{i}))\to\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}_{\mathrm{Roe}}(L)).$
  2. (2)
    The following diagram commutes:
[…]
Proof.
(1) follows from Proposition 3.19 as both $L_{1}$ and $L_{2}$ are coarsely dense in $\mathbb{R}^{d}$, hence in $L_{0}$. In particular, we have a direct sum decomposition of Hilbert spaces
$\mathcal{H}_{0}=\mathcal{H}_{1}\oplus\mathcal{H}_{2},$
where $\mathcal{H}_{i}$ is identified with the image of the isometry $V_{i}\colon\mathcal{H}_{i}\hookrightarrow\mathcal{H}_{0}$.
Now we prove (2). Let $i\in\{0,1,2\}$. The position spectral triples $\xi^{\mathrm{Roe}}_{i}$ for $i\in\{0,1,2\}$ are given by the position operators $\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}^{i}_{j}\otimes\gamma^{j}$, where ${\scriptstyle\mathsf{X}}_{j}^{i}$ is the $j$-th position operator on $\mathcal{H}_{i}$:
${\scriptstyle\mathsf{X}}^{i}_{j}f(x)\mathrel{\vcentcolon=}x_{j}f(x),\quad x\in L_{i};$
The operators ${\scriptstyle\mathsf{X}}^{0}_{j}$ is diagonal for the decomposition $\mathcal{H}_{0}=\mathcal{H}_{1}\oplus\mathcal{H}_{2}$. Its restriction to $\mathcal{H}_{i}$ is ${\scriptstyle\mathsf{X}}^{i}_{j}$. Therefore, we have
${\scriptstyle\mathsf{X}}^{0}_{j}=V_{1}{\scriptstyle\mathsf{X}}_{j}^{1}V_{1}^{*}\oplus V_{2}{\scriptstyle\mathsf{X}}_{j}^{2}V_{2}^{*}=\operatorname{Ad}_{V_{1}}({\scriptstyle\mathsf{X}}_{j}^{1})\oplus\operatorname{Ad}_{V_{2}}({\scriptstyle\mathsf{X}}_{j}^{2}),$
and hence
$\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}^{0}_{j}\otimes\gamma^{j}=\operatorname{Ad}_{V_{1}}\left(\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}^{1}\otimes\gamma^{j}\right)\oplus\operatorname{Ad}_{V_{2}}\left(\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}^{2}\otimes\gamma^{j}\right).$
Therefore, the *-homomorphism $\operatorname{Ad}_{V_{1}}$ (resp. $\operatorname{Ad}_{V_{2}}$) pulls back the K-homology class represented by the position spectral triple $\xi_{0}^{\mathrm{Roe}}$ to the class represented by $\xi_{1}^{\mathrm{Roe}}$ (resp. $\xi_{2}^{\mathrm{Roe}}$ ). So the left (resp. right) triangle commutes by Lemma 2.13. ∎
Theorem 3.32.
For any Delone set $\Lambda$, the position spectral triple $\xi^{\mathrm{Roe}}_{\Lambda}$ induces an isomorphism
$\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda))\xrightarrow{\sim}\operatorname{KK}(\mathrm{C}\ell_{d,d},\mathbb{R})\simeq\mathbb{Z}.$
Proof.
Set $L_{1}=\mathbb{Z}^{d}$ and $L_{2}=\Lambda$. Then the position spectral triple $\xi^{\mathrm{Roe}}_{1}=\xi^{\mathrm{Roe}}_{\mathbb{Z}^{d}}$ induces an isomorphism
$\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}))\xrightarrow{\sim}\operatorname{KK}(\mathrm{C}\ell_{d,d},\mathbb{R})\simeq\mathbb{Z}$
by Lemma 3.26. The isometry $V_{1}$ induces an isomorphism between the corresponding Kasparov groups by Lemma 3.30. Thus the commutativity of the left triangle in (3.31) implies that the map
$\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}\cup\Lambda))\to\mathbb{Z}$
given by the position spectral triple $\xi^{\mathrm{Roe}}_{\mathbb{Z}^{d}\cup\Lambda}$, must be an isomorphism. Thus the position spectral triple $\xi_{2}^{\mathrm{Roe}}=\xi_{\Lambda}^{\mathrm{Roe}}$ must be an isomorphism as well, as it is a composition of isomorphisms
$\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda))\underset{\sim}{\xrightarrow{\iota_{1*}}}\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}\cup\Lambda))\underset{\sim}{\xrightarrow{\xi_{0}^{\mathrm{Roe}}}}\mathbb{Z}.\qed$

### v2

@@ 3.2.2. Position spectral triple generates $\operatorname{KO}_{*}(\mathbb{R})$
The proof comes from [Ewert-Meyer:Coarse_geometry]*Theorem 7, Corollary 5, which we sketch here.
First let $n=d$ and $\iota\colon\mathrm{C}^{*}(\mathbb{Z}^{d})\hookrightarrow\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})$ be the inclusion map induced by the obvious isometry $\ell^{2}(\mathbb{Z}^{d})\hookrightarrow\ell^{2}(\mathbb{Z}^{d})\otimes\mathcal{K}$. Then
$\xi_{\mathbb{Z}^{d},1}=\iota^{*}\xi_{\mathbb{Z}^{d}}^{\mathrm{Roe}},$
that is, $\xi_{\mathbb{Z}^{d},1}$ is the pullback of the position spectral triple over $\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})_{\mathbb{R}}$. Thus the following diagram commutes by functorality of the Kasparov product (cf. Lemma˜2.10):
$\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}})$$\operatorname{KK}(\mathrm{C}\ell_{d,d},\mathbb{R})\simeq\mathbb{Z}$$\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})_{\mathbb{R}})$$\scriptstyle{\times_{\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}}[\xi_{\mathbb{Z}^{d},1}]}$$\scriptstyle{\times_{A}[\iota]}$$\scriptstyle{\times_{\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})_{\mathbb{R}}}[\xi^{\mathrm{Roe}}_{\mathbb{Z}^{d}}]}$
and the horizontal arrow is a surjective group homomorphism by Theorem˜2.15. Thus the bottom right arrow is a surjective group homomorphism $\mathbb{Z}\to\mathbb{Z}$, hence an isomorphism.
The isomorphisms of abelian groups $\operatorname{KO}_{n}(\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})_{\mathbb{R}})\simeq\operatorname{KO}_{n-d}(\mathbb{R})$ give a give an isomorphism of $\operatorname{KO}_{*}(\mathbb{R})$-modules $\operatorname{KO}_{*}(\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})_{\mathbb{R}})\simeq\operatorname{KO}_{*-d}(\mathbb{R})$ (cf. the discussion after [Ewert-Meyer:Coarse_geometry]*Proposition 8). Functorality of the Kasparov product implies that $\xi_{\mathbb{Z}^{d}}^{\mathrm{Roe}}$ induces a $\operatorname{KO}_{*}(\mathbb{R})$-module homomorphism $\operatorname{KO}_{*}(\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})_{\mathbb{R}})\to\operatorname{KO}_{*-d}(\mathbb{R})$, both sides being isomorphic to free $\operatorname{KO}_{*}(\mathbb{R})$-modules with a generator in degree d-d. The commutative diagram above shows that $\xi_{\mathbb{Z}^{d}}^{\mathrm{Roe}}]$ maps such a generator to a generator, hence a $\operatorname{KO}_{*}(\mathbb{R})$-module isomorphism. In particular, it follows that the induced maps
$\xi_{\mathbb{Z}^{d}}^{\mathrm{Roe}}\colon\underbrace{\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})_{\mathbb{R}})}_{\simeq\operatorname{KO}_{n}(\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})_{\mathbb{R}})}\to\underbrace{\operatorname{KK}(\mathrm{C}\ell_{n,d},\mathbb{R})}_{\simeq\operatorname{KO}_{n-d}(\mathbb{R})}$
are isomorphisms of abelian groups for all $n$. ∎
Lemma 3.31.
Let $L_{1},L_{2}\subseteq\mathbb{R}^{d}$ be Delone sets such that $L_{1}\cap L_{2}=\emptyset$. Let $L\mathrel{\vcentcolon=}L_{1}\sqcup L_{2}$. For $i\in\{1,2\}$, we write:
  • $\mathcal{H}_{i}\mathrel{\vcentcolon=}\ell^{2}(L_{i},\mathcal{K})$ for the standard ample $L_{i}$-module, and $\mathcal{H}\mathrel{\vcentcolon=}\ell^{2}(L_{1}\sqcup L_{2},\mathcal{K})$ for the standard ample $L$-module;
  • $\xi^{\mathrm{Roe}}_{L_{i}}$ and $\xi^{\mathrm{Roe}}_{L}$ for the position spectral triples over $\mathrm{C}^{*}_{\mathrm{Roe}}(L_{i})$ and $\mathrm{C}^{*}_{\mathrm{Roe}}(L)$ as in (3.28);
  • $\iota_{i}\colon L_{i}\hookrightarrow L$ for the isometric embeddings;
  • $V_{i}\colon\mathcal{H}_{i}\hookrightarrow\mathcal{H}_{0}$ for the isometries induced by $\iota_{i}$, that is, $V_{i}\lvert x\rangle\mathrel{\vcentcolon=}\lvert\iota_{i}(x)\rangle$ for all $x\in L_{i}$.
Then the following hold:
  1. (1)
    $V_{i}^{\prime}s$ are covering isometries for $\iota_{i}$. Hence $\operatorname{Ad}_{V_{i}}\colon T\mapsto V_{i}TV_{i}^{*}$ maps $\mathrm{C}^{*}_{\mathrm{Roe}}(L_{i})$ into $\mathrm{C}^{*}_{\mathrm{Roe}}(L)$, and induces an isomorphism
    $\iota_{i*}\colon\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(L_{i}))\to\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(L)).$
  2. (2)
    The following diagram commutes (all arrows are given by taking the Kasparov product):
    (3.32) $\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(L_{1}))$$\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(L))$$\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(L_{2}))$$\operatorname{KK}(\mathrm{C}\ell_{n,d},\mathbb{R})$$\scriptstyle{\iota_{1*}}\scriptstyle{\sim}$$\scriptstyle{\xi^{\mathrm{Roe}}_{L_{1}}}$$\scriptstyle{\xi^{\mathrm{Roe}}_{L}}$$\scriptstyle{\iota_{2*}}\scriptstyle{\sim}$$\scriptstyle{\xi^{\mathrm{Roe}}_{L_{2}}}$
Proof.
(1) follows from Proposition˜3.22 as both $L_{1}$ and $L_{2}$ are coarsely dense in $L$. Since $L_{1}\cap L_{2}=\emptyset$ and $L=L_{1}\sqcup L_{2}$, it follows that
$\mathcal{H}_{1}\oplus\mathcal{H}_{2}\xrightarrow{\sim}\mathcal{H},\quad(\phi_{1},\phi_{2})\mapsto V_{1}\phi_{1}+V_{2}\phi_{2}.$
is a unitary isomorphism of Hilbert spaces.
Let ${\scriptstyle\mathsf{X}}^{i}_{j}$ be the $j$-th position operator on $\ell^{2}(L_{i})$, that is,
${\scriptstyle\mathsf{X}}_{j}^{i}f(x)\mathrel{\vcentcolon=}x_{j}f(x),\quad\phi\in\mathrm{C}_{\mathrm{c}}(L_{i}),\;x\in L_{i}.$
Also write ${\scriptstyle\mathsf{X}}_{j}$ for the $j$-th position operator on $\ell^{2}(L)$. Then ${\scriptstyle\mathsf{X}}_{j}$ restricts to the $j$-th position operator on $\ell^{2}(L_{i})$. That is, we have
${\scriptstyle\mathsf{X}}_{j}=V_{1}{\scriptstyle\mathsf{X}}_{j}^{1}V_{1}^{*}\oplus V_{2}{\scriptstyle\mathsf{X}}_{j}^{2}V_{2}^{*}=\operatorname{Ad}_{V_{1}}({\scriptstyle\mathsf{X}}_{j}^{1})\oplus\operatorname{Ad}_{V_{2}}({\scriptstyle\mathsf{X}}_{j}^{2}),$
and hence
$\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\operatorname{id}_{\mathcal{K}}\otimes\gamma^{j}=\operatorname{Ad}_{V_{1}}\left(\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}^{1}\otimes\operatorname{id}_{\mathcal{K}}\otimes\gamma^{j}\right)\oplus\operatorname{Ad}_{V_{2}}\left(\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}^{2}\otimes\operatorname{id}_{\mathcal{K}}\otimes\gamma^{j}\right).$
Therefore, for $i\in\{1,2\}$, the *-homomorphism $\operatorname{Ad}_{V_{j}}$ pulls back the class $\xi_{L}^{\mathrm{Roe}}]\in\operatorname{KK}(\mathrm{C}^{*}_{\mathrm{Roe}}(L)\otimes\mathrm{C}\ell_{0,d},\mathbb{R})$ to the class $\xi_{L_{i}}^{\mathrm{Roe}}]\in\operatorname{KK}(\mathrm{C}^{*}_{\mathrm{Roe}}(L_{i})\otimes\mathrm{C}\ell_{0,d},\mathbb{R})$. So both the left and the right triangle commutes by Lemma˜2.10. ∎
Theorem 3.33.
For any Delone set $\Lambda$, the position spectral triple $\xi^{\mathrm{Roe}}_{\Lambda}$ induces, for each $n$, an isomorphism
$\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda))\xrightarrow{\sim}\operatorname{KK}(\mathrm{C}\ell_{n,d},\mathbb{R}).$
Proof.
The first step is to replace $\Lambda$ by its translated copy that is disjoint from $\mathbb{Z}^{d}$. To be precise, we note that the set
$B\mathrel{\vcentcolon=}\{x-y\mid x\in\Lambda,y\in\mathbb{Z}^{d}\}$
is countable, so the set $\mathbb{R}^{d}\setminus B$ is non-empty. Choose any $v\in\mathbb{R}^{d}\setminus B$, then the Delone set $\Lambda^{\prime}\mathrel{\vcentcolon=}\Lambda-v$ is disjoint from $\mathbb{Z}^{d}$. The translation map
$f\colon\Lambda\to\Lambda^{\prime},\quad x\mapsto x-v$
induces a unitary isomorphism of Hilbert spaces $V\colon\ell^{2}(\Lambda)\xrightarrow{\sim}\ell^{2}(\Lambda^{\prime})$ that covers $f$. Hence $V$ gives a *-isomorphism $\operatorname{Ad}_{V}\colon\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)\xrightarrow{\sim}\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda^{\prime})$, which pulls back the position spectral triple $\xi_{\Lambda^{\prime}}^{\mathrm{Roe}}$ to the position spectral triple $\xi_{\Lambda}^{\mathrm{Roe}}$.
As a consequence, we may assume without loss of generality that $\mathbb{Z}^{d}\cap\Lambda=\emptyset$. Set $L_{1}\mathrel{\vcentcolon=}\mathbb{Z}^{d}$ and $L_{2}\mathrel{\vcentcolon=}\Lambda$. Then the position spectral triple $\xi^{\mathrm{Roe}}_{L_{1}}=\xi^{\mathrm{Roe}}_{\mathbb{Z}^{d}}$ induces an isomorphism
$\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}))\xrightarrow{\sim}\operatorname{KK}(\mathrm{C}\ell_{n,d},\mathbb{R})$
by Lemma˜3.29. The isometry $V_{1}$ induces an isomorphism between the corresponding Kasparov groups by Lemma˜3.31. Thus the commutativity of the left triangle in (3.32) implies that the map
$\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}\sqcup\Lambda))\to\operatorname{KK}(\mathrm{C}\ell_{n,d},\mathbb{R})$
given by the position spectral triple $\xi^{\mathrm{Roe}}_{\mathbb{Z}^{d}\sqcup\Lambda}$, must be an isomorphism. Thus the position spectral triple $\xi_{L_{2}}^{\mathrm{Roe}}=\xi_{\Lambda}^{\mathrm{Roe}}$ must be an isomorphism as well, as it is a composition of isomorphisms
$\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda))\underset{\sim}{\xrightarrow{\iota_{1*}}}\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}\sqcup\Lambda))\underset{\sim}{\xrightarrow{\xi_{L}^{\mathrm{Roe}}}}\operatorname{KK}(\mathrm{C}\ell_{n,d},\mathbb{R}).\qed$
3.2.3. Remarks on uniform Roe C*-algebras
Let $\Lambda$ be a discrete metric space. An operator $T\in\mathbb{B}(\ell^{2}(\Lambda))$ is called controlled, if there exists $R>0$ such that $T_{x,y}\mathrel{\vcentcolon=}\langle x\,|T|\,\mathopen{}y\rangle=0$ whenever $x,y\in\Lambda$ satisfy $\mathrm{d}(x,y)>R$. The uniform Roe C*-algebra of $\Lambda$, denoted by $\mathrm{C}^{*}_{\mathrm{u,Roe}}(\Lambda)$, is the C*-algebra generated by all controlled operators on $\ell^{2}(\Lambda)$.
Starting from Kubota [Kubota:Controlled_topological_phases], uniform Roe C*-algebras have also been employed as the observable C*-algebras of aperiodic topological insulators. As opposed to Roe C*-algebras, uniform Roe C*-algebras have much more involved K-theory. For example, it is known that $\operatorname{K}_{0}(\mathrm{C}^{*}_{\mathrm{u,Roe}}(\mathbb{Z}))$ is uncountable for $d\geq 1$ (cf. [Spakula:K-theory_uniform_Roe_algebras]*Example II.3.4).
We briefly remark here that the uniform Roe C*-algebra is also quite “universal” as it contains all “tight-binding” observable C*-algebras, in the following sense:
Proposition 3.34.
Assume that there is a real spectral triple of the form
(3.35) $\zeta_{\Lambda}\mathrel{\vcentcolon=}\left(\mathcal{A}\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\Lambda)\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad{\scriptstyle\mathsf{X}}\mathrel{\vcentcolon=}\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right),$
where:
  • AA is a real C*-algebra, which carries a *-representation $\varphi\colon A\to\mathbb{B}(\ell^{2}(\Lambda))$; $\mathcal{A}\subseteq A$ is a dense *-subalgebra;
  • ${\scriptstyle\mathsf{X}}_{j}$ is the $j$-th position operator on $\ell^{2}(\Lambda)$, given by
    $({\scriptstyle\mathsf{X}}_{j}\phi)(x)\mathrel{\vcentcolon=}x_{j}\phi(x),\qquad\phi\in\mathrm{C}_{\mathrm{c}}(\Lambda),\;x=(x_{1},\dots,x_{d})\in\Lambda\subseteq\mathbb{R}^{d};$
Then $\varphi(A)$ is contained in $\mathrm{C}^{*}_{\mathrm{u,Roe}}(\Lambda)$.
We note that the auxiliary Hilbert space $\mathcal{K}$ is now absent, as opposed to the definition of a position spectral triple (cf. Definition˜2.11). Thus $A$ is merely represented on the “tight-binding” Hilbert space $\ell^{2}(\Lambda)$ instead of $\ell^{2}(\Lambda)\otimes\mathcal{K}$. Replacing $A$ by $A\otimes\mathbb{M}_{N}(\mathbb{R})$ replaces $\ell^{2}(\Lambda)$ by $\ell^{2}(\Lambda,\mathbb{R}^{N})=\ell^{2}(\Lambda)\otimes\mathbb{R}^{N}$.
Proof.
It follows from Lemma˜3.25 and the proof of Theorem˜3.26 that given the spectral triple $\zeta_{\Lambda}$, then every $a\in\mathcal{A}$ is represented as a controlled operator on $\ell^{2}(\Lambda)$. Thus $\varphi(\mathcal{A})\subseteq\mathrm{C}^{*}_{\mathrm{u,Roe}}(\Lambda)$ and hence $\varphi(A)\subseteq\mathrm{C}^{*}_{\mathrm{u,Roe}}(\Lambda)$. ∎
Thus tight-binding models of topological insulators, e.g. the groupoid model $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$, is already contained in the uniform Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{u,Roe}}(\Lambda)$, and we have inclusions of C*-algebras and isometric embeddings of Hilbert spaces which they act on:
@@ 3.2.3. Remarks on uniform Roe C*-algebras
As argued in [Ewert-Meyer:Coarse_geometry]*Section 1, the occurrence of uniform Roe C*-algebras might be considered as an artefact of the tight-binding approximation.